# Studying propagating turbulent structures in the near wake of a sphere using Hilbert proper orthogonal decomposition

### v1

@@ 2.2 Hilbert Proper Orthogonal Decomposition
HPOD embeds the phase information in the modes by replacing $\boldsymbol{f}(\boldsymbol{x},t)$ with its analytic signal [32]
@@
where $\mathcal{H}$ denotes the Hilbert transform of $\boldsymbol{f}(\boldsymbol{x},t)$ in time
[…]
The Hilbert transform is equivalent to a $\mp 90^{\circ}$ phase shift [36] of the fundamental frequencies of $\boldsymbol{f}(\boldsymbol{x},t)$ in Fourier space
@@
For time-resolved data, the Hilbert transform can be performed in time to determine the analytic signal $\boldsymbol{f}^{a}(\boldsymbol{x},t)$. HPOD is performed on time-resolved data by replacing $\boldsymbol{f}(\boldsymbol{x},t)$ with its analytic signal in equations (4) to (6). However, where the coherent structures propagate downstream, the Hilbert transform can be performed in the streamwise direction, as propagation in this direction serves as an analogy for a change in time, to apply this method to data that is not time-resolved [33, 34, 35]
[…]
In this case, the analytic signal of $\boldsymbol{f}(\boldsymbol{x},t)$ is defined using $\mathcal{H}_{x}$
@@
and substituted for $\boldsymbol{f}(\boldsymbol{x},t)$ in equations (7) to (9). The resulting eigenvalues $\lambda^{a}_{k}$ are the contribution of the $k^{th}$ mode to the total TKE of the analytic signal, and are real-valued. The spatial modes $\psi^{a}_{k}(\boldsymbol{x})$ are complex, with the imaginary part of $\psi^{a}_{k}(\boldsymbol{x})$ representing a phase shift of its real part by $\mp 90^{\circ}$.
The instantaneous amplitude and phase angle of the $k^{th}$ mode are defined by the real $\Re$ and imaginary $\Im$ parts of $a^{a}_{k}(t)$ as
@@
respectively. The instantaneous amplitude denotes the strength of the mode at each instant in time, and the instantaneous phase angle denotes the mode shape, which oscillates between the real and imaginary parts of $\psi_{k}^{a}(\boldsymbol{x})$ at $\phi_{k}^{a}(t)$ equal to $0^{\circ}$ and $\mp 90^{\circ}$, respectively.
The analytic signal can be reconstructed using $\hat{K}$ modes by
@@
the real part of which is the reduced-order model of $\boldsymbol{f}(\boldsymbol{x},t)$ including the modes which contribute the most TKE to the analytic signal. This reconstruction favours the propagating components of $\boldsymbol{f}(\boldsymbol{x},t)$, but introduces artefacts due to the Hilbert transform.
@@ 2.3 Hilbert Transform of POD Modes
Although the POD modes do not contain temporal information, the streamwise Hilbert transform can be performed directly on each of the modes. Thus, the analytic signal of the POD mode $\psi_{k}(\boldsymbol{x})$ can be defined as
@@
$\sum_{k=1}^{K}a_{k}(t)\Big{(}(\psi_{k})^{a}(\boldsymbol{x})\Big{)}=\sum_{k=1}^{K}a_{k}(t)\psi_{k}(\boldsymbol{x})+a_{k}(t)i\mathcal{H}_{x}\big{[}\psi_{k}(\boldsymbol{x})\big{]}.$ (25)
As the Hilbert transform is a linear operator, this can be rewritten as
[…]
@@
Therefore, applying the transform to each of the POD modes of $\boldsymbol{f}(\boldsymbol{x},t)$ individually and using the resulting analytic signals in place of the original modes allows for the reconstruction of $\boldsymbol{f}^{a}(\boldsymbol{x},t)$. However, the analytic signal of the POD modes are not equal to the HPOD modes due to the filtering of the Hilbert transform being introduced after the decomposition of $\boldsymbol{f}(\boldsymbol{x},t)$. While this method is not equivalent to HPOD, where the propagation of structures is embedded in the modes, it is a computationally efficient alternative for matching pairs of POD modes which represent the same propagating structure.
@@ 3.1 Experimental Facility
Experiments were performed in a vertical water tunnel, the test section of which is 1.5 m long with a cross-section of 0.25 m $\times$ 0.25 m, shown in figure 1[37, 38, 39]. The settling chamber is connected to the top of the test section by a 16:1 contraction, which includes a honeycomb and screens to reduce the turbulence level and scale in the tunnel test section. Water is pumped from the plenum chamber below the test section up to the settling chamber such that the flow through the test section is downwards. The walls of the test section are made of 15 mm thick clear acrylic. A removable panel is located on one wall for access to the test section, and is set flush to the internal side of the wall to minimise asymmetry in the tunnel. Seeding and filtration of the tunnel is performed via an auxiliary circuit, which is isolated during experiments. A sting of diameter 9.3 mm is mounted to a crossbeam of thin aerofoil cross-sections above the test section, and tapers in diameter over the last 100 mm down to 4.65 mm in diameter and ends in a threaded rod which is used to mount a sphere in the test section. A close-up view of the sphere mounting structure is shown in figure 1b.
A sphere with a diameter of $D=40$ mm was manufactured using a Phrozen Sonic Mighty 8k 3-D printer, which has a 35 µm layer height and a planar resolution of 22 µm $\times$ 22 µm, as shown in figure 1c. The sphere has a shell thickness of 2.5 mm and an external boss at the top for mounting to a threaded rod on the end of the sting. The boss begins parallel to the sting and smoothly transitions to the sphere surface. The sphere is supported internally by a conical structure, which features slits to prevent uncured resin from being trapped in the cavities during printing.
Figure 1: (a) Vertical water tunnel facility, and (b) close-up of the mounting structure and sphere, and (c) 3-D printed sphere design [39].
@@ 3.2 Experimental Method
Measurements of the flow in the wake of the sphere were taken with a freestream velocity of $U_{\infty}=200$ m/s, resulting in a Reynolds number of $Re_{D}=7780$. Under these conditions, the background turbulence is 2%, and the mean convective time of the flow
@@
is 0.2 seconds [39]. The flow was seeded with 11 µm glass spheres, and the single-exposed particle image pairs were recorded using a PCO Panda 26 DS camera with a 5120 $\times$ 5120 px array of 2.5 µm $\times$ 2.5 µm pixels. A Zeiss lens with a fixed 50 mm focal length was used to capture a 190 mm $\times$ 190 mm field of view, resulting in a spatial resolution is 37 µm/px. The field of view began just behind the trailing edge of the sphere, to avoid reflections off the surface, and extended to 5.2 diameters downstream of the sphere. The field of view was aligned with the centreline of the sphere, and extended to 2.3 diameters either side of the centreline in the transverse direction. The f-stop was set to 2.8, resulting in a 5.0 mm depth of field. The diffraction limited minimum image diameter and expected particle image diameter for an 11 µm particle at these setting are 3.9 µm and 1.6 px, respectively.
The particle images were illuminated using 7 ns pulses from a pair of 400mJ Spectra Physics Nd:YAG lasers. The beams of these lasers were aligned co-linearly before being shaped into a sheet aligned to the centre of the tunnel and the corresponding meridian of the sphere. The thickness of the laser sheet throughout the field of view was 300 µm. Pairs of single-exposed images where captured with a separation of $\Delta t=1$ ms at an acquisition rate of 1 Hz ($\Delta T=1$ s), using a BeagleBone Black pulse generator [40] to control the timing and synchronisation of the camera and laser pulses.
@@ 3.3 Data Processing
The mean background images of the first and second exposures were subtracted from the corresponding single-exposed particle image pairs. The instantaneous streamwise and transverse velocities, $u(x,y,t)$ and $v(x,y,t)$, were determined from the particle image pairs using two-component-two-dimensional (2C-2D) multigrid/multipass cross-correlation digital (MCCD) PIV [41, 42]. A final interrogation window of 64 px $\times$ 32 px on a grid with spacing of 32 px $\times$ 16 px, in $x$ and $y$, respectively, was used. This aspect ratio was chosen to capture the higher streamwise velocity component and the higher velocity gradients in the transverse direction, while still being sufficiently large for the interrogation window to contain approximately 10 particles common to both images of each pair. The quality of the cross-correlation was increased through the application of an 8-point Hart filter. A maximum displacement limit of 20 px and normalised local median filter [43], with a threshold of 2 standard errors, was applied to the raw displacement vectors. The validation criteria resulted in a rejection of 2.3% of vectors, which were replaced by an interpolation of a 2nd-order polynomial fit of their 13 nearest neighbours [42]. A total of $N_{T}=12,600$ two-component velocity vector fields, each containing 50,000+ vectors, were determined. The instantaneous velocity fluctuations $u^{\prime}(x,y,t)$ and $v^{\prime}(x,y,t)$ were calculated by subtracting the mean velocity field, $\overline{u}(x,y)$ and $\overline{v}(x,y)$, from the instantaneous velocity field, $u(x,y,t)$ and $v(x,y,t)$.
The 95% confidence interval of the 2C-2D MCCDPIV algorithm used is 0.06 pixels [42], which corresponds to a standard error of $\sigma_{\varepsilon_{u}}=0.03$ pixels. The standard error of the instantaneous velocities is thus $1.1\times 10^{-3}$ m/s. The uncertainty of the mean velocity arising from the random nature of the turbulent fluctuations can be estimated by including the mean streamwise Reynolds stress [44, 45]
@@
where $\sigma_{\varepsilon_{u}}$ is the standard error of the instantaneous streamwise velocity measurements. As the maximum of both $\overline{u^{\prime}u^{\prime}}$ and $\overline{v^{\prime}v^{\prime}}$ are of the order of $0.05U_{\infty}^{2}$ [39], a conservative estimate of the uncertainty in the mean velocity components is $4.0\times 10^{-4}$ m/s. The uncertainty in the fluctuations is thus
@@
Freestream velocity $U_{\infty}$ 200 mm/s --
Sphere diameter $D$ 40 mm --
Reynolds number $Re_{D}$ 7780 --
Convective time $t_{c}$ 0.2 s --
@@
Depth of field -- 5.0 mm $0.125D$
Laser sheet thickness -- 0.3 mm $0.0075D$
Diffraction limited particle diameter -- 3.9 µm
Spatial resolution $SR$ 37 µm --
Grid spacing -- 32 px $\times$ 16 px $0.03D\times 0.015D$
Interrogation window -- 64 px $\times$ 32 px $0.06D\times 0.06D$
Field of view -- 190 mm $\times$ 190 mm $4.7D\times 4.7D$
Number of 2C-2D velocity fields $N_{T}$ 12,600 --
Velocity vectors per field $N_{S}$  50,000+ --
Velocity vector validation rate -- 97.7%
@@ 4.1 POD and HPOD Modes
POD was performed by arranging the turbulent fluctuations into the matrix
@@
where $N_{S}$ is the number of velocity vectors per velocity field and $\boldsymbol{x}_{n}={(x,y)}_{n}$ is the spatial coordinates of the $n^{th}$ vector, which has fluctuating components denoted by $u^{\prime}$ and $v^{\prime}$ in the streamwise and transverse directions, respectively. As $N_{S}>N_{T}$, the method of snapshots was used to perform the POD, with the two-point correlation of $\boldsymbol{X}$ in time given by the matrix form of equation (7)
@@
yielded the eigenvectors $\boldsymbol{\Phi}$ and eigenvalues $\lambda$. The temporal coefficients were determined from $\boldsymbol{\Phi}$
@@
and the spatial modes were given by the projection of $\boldsymbol{X}$ onto $\boldsymbol{\Phi}$
@@
The analytic signal of $\boldsymbol{X}$
$\boldsymbol{X}^{a}=\boldsymbol{X}+i\mathcal{H}_{x}[\boldsymbol{X}],$ (35)
was determined using equations (18) and (19), and HPOD was performed by substituting $\boldsymbol{X}^{a}$ for $\boldsymbol{X}$ in equations (31) to (34). The resulting eigenvalues, temporal coefficients and spatial modes are denoted by $\lambda^{a}_{k}$, $a_{k}^{a}(t)$ and $\psi_{k}^{a}(\boldsymbol{x})$, respectively.
The contribution of the leading HPOD modes to the TKE of $\boldsymbol{X}^{a}$ is significantly larger than the contribution of the leading POD modes to the TKE of $\boldsymbol{X}$. The first, second and third modes of each decomposition contribute 7.70%, 4.53% and 3.70%, and 4.8%, 3.39% and 2.55%, respectively. However, the difference in the contribution of the higher order modes to the TKE is less significant, as shown in figure 2a. The result is that the cumulative contribution to TKE of a given number of the HPOD modes is larger than that of the equivalent number of POD modes, as shown in figure 2b and table 2. However, this is not a one-to-one comparison, as the total TKE of the analytic signal $\boldsymbol{X}^{a}$ is larger than that of the data $\boldsymbol{X}$, at 571,871 m2/s2 and 300,542 m2/s2, respectively. This is due to the imaginary component of $\boldsymbol{X}^{a}$, which is not present in $\boldsymbol{X}$.
@@
[…]
@@
The sixth and eighth POD modes, shown in figures 3f and 3h, respectively, are similar in structure to the first and second POD modes but with a shorter wavelength than those modes. The seventh POD mode superficially resembles the fourth mode in its transverse component, shown in figure 3g(ii). However, the streamwise component, shown in figure 3g(i), does not exhibit the streamwise periodicity of the other leading modes. While this mode may represent an asymmetric aspect of the structure represented by the forth and fifth modes, it is less clearly related to these modes than the third mode appears to be to the first and second modes. The ninth and tenth POD modes respectively exhibit V-shaped structures in their streamwise components, shown in figures 3i(i) and 3j(i), respectively, similar to the fourth and fifth modes for $x/D>2.5$, which increase in wavelength further downstream. These modes differ in shape from the lower-order modes upstream of $x/D=2$, where the structures have an aspect ratio closer to unity than those in the higher-order modes. The transverse components of the ninth and tenth modes, shown in figures 3i(ii) and 3j(ii), respectively, resemble those of the fourth and fifth modes with a short wavelength. The wavelength of these modes also increase in the downstream direction.
@@
Figure 4: (i) Streamwise and (ii) transverse components of the first four HPOD modes.

### v2

@@ 2.2 Hilbert Proper Orthogonal Decomposition
HPOD incorporates phase information into the modes by replacing $\boldsymbol{f}(\boldsymbol{x},t)$ with its analytic signal [16]
@@
where $\mathcal{H}$ denotes the Hilbert transform of $\boldsymbol{f}(\boldsymbol{x},t)$ with respect to time
$\mathcal{H}\big[\boldsymbol{f}(\boldsymbol{x},t)\big]=\frac{1}{\pi}\int_{-\infty}^{\infty}\frac{\boldsymbol{f}(\boldsymbol{x},\tau)}{t-\tau}\,d\tau.$ (15)
The Hilbert transform is equivalent to a $\mp\pi/2$ phase shift of the fundamental frequency components of $\boldsymbol{f}(\boldsymbol{x},t)$ in Fourier space [40]
@@
For time-resolved data, the Hilbert transform is applied with respect to time to compute the analytic signal $\boldsymbol{f}^{a}(\boldsymbol{x},t)$. HPOD is implemented on time-resolved data by substituting $\boldsymbol{f}(\boldsymbol{x},t)$ in equations (4) to (6) with its analytic signal. For flows where coherent structures propagate downstream, the Hilbert transform can alternatively be applied in the streamwise direction
$\mathcal{H}_{x}\big[\boldsymbol{f}(\boldsymbol{x},t)\big]=\frac{1}{\pi}\int_{-\infty}^{\infty}\frac{\boldsymbol{f}\big((\xi,\ldots),t\big)}{x-\xi}\,d\xi.$ (18)
This approach uses streamwise propagation as an analogue for temporal evolution, allowing the method to be applied to data that are not time-resolved [27, 28, 20].
In this case, the analytic signal of $\boldsymbol{f}(\boldsymbol{x},t)$ is defined using the streamwise Hilbert transform $\mathcal{H}_{x}$
@@
and is substituted for $\boldsymbol{f}(\boldsymbol{x},t)$ in equations (7) to (9). The resulting eigenvalues $\lambda^{a}_{k}$ represent the contribution of the $k^{th}$ mode to the total TKE of the analytic signal and are real-valued. The spatial modes $\psi^{a}_{k}(\boldsymbol{x})$ are complex, with the imaginary component of $\psi^{a}_{k}(\boldsymbol{x})$ corresponding to a $\mp\pi/2$ phase shift of the real component.
The instantaneous amplitude and phase angle of the $k^{th}$ mode are defined by the real $\Re$ and imaginary $\Im$ components of $a^{a}_{k}(t)$ as
@@
respectively. The instantaneous amplitude quantifies the strength of the mode at each instant, while the instantaneous phase angle characterises the mode shape, which oscillates between the real and imaginary components of $\psi^{a}_{k}(\boldsymbol{x})$ at $\phi^{a}_{k}(t)=0^{\circ}$ and $\phi^{a}_{k}(t)=\mp\pi/2$, respectively.
The analytic signal can be reconstructed using $\hat{K}$ modes as
@@
where the real part is the reduced-order model of $\boldsymbol{f}(\boldsymbol{x},t)$ including the modes which contribute most to the TKE of the analytic signal. This reconstruction favours the propagating components of $\boldsymbol{f}(\boldsymbol{x},t)$, but introduces spectral leakage into the reconstruction due to the assumption of periodicity inherent to Hilbert transform.
@@ 2.3 Hilbert Transform of POD Modes
Although POD modes do not inherently contain temporal information, the streamwise Hilbert transform can be applied directly to each mode. Thus, the analytic signal of the POD mode $\psi_{k}(\boldsymbol{x})$ is defined as
@@
$\sum_{k=1}^{K}a_{k}(t)\Big({(\psi_{k})}^{a}(\boldsymbol{x})\Big)=\sum_{k=1}^{K}a_{k}(t)\psi_{k}(\boldsymbol{x})+\sum_{k=1}^{K}a_{k}(t)i\mathcal{H}_{x}\big[\psi_{k}(\boldsymbol{x})\big].$ (25)
Since the Hilbert transform is a linear operator, the expression can be rewritten as
$\sum_{k=1}^{K}a_{k}(t)\Big({(\psi_{k})}^{a}(\boldsymbol{x})\Big)$ $=\sum_{k=1}^{K}a_{k}(t)\psi_{k}(\boldsymbol{x})+i\mathcal{H}_{x}\Big[\sum_{k=1}^{K}a_{k}(t)\psi_{k}(\boldsymbol{x})\Big]$ (26)
$=\boldsymbol{f}(\boldsymbol{x},t)+i\mathcal{H}_{x}[\boldsymbol{f}(\boldsymbol{x},t)]$
$=\boldsymbol{f}^{a}_{x}(\boldsymbol{x},t),$
@@
Therefore, applying the Hilbert transform to each POD mode of $\boldsymbol{f}(\boldsymbol{x},t)$ individually and substituting the resulting analytic signals for the original modes enables the reconstruction of $\boldsymbol{f}^{a}(\boldsymbol{x},t)$. However, the analytic signals of the POD modes are not equal to the HPOD modes, as the decomposition of the analytic signal of $\boldsymbol{f}(\boldsymbol{x},t)$ ranks HPOD modes by their energy contribution to $\boldsymbol{f}^{a}(\boldsymbol{x},t)$ rather than to $\boldsymbol{f}(\boldsymbol{x},t)$. Consequently, leading POD modes may contain non-propagating features that are relegated to higher-order HPOD modes. While this approach is not equivalent to HPOD, the tendency for propagating structures to be decomposed into paired POD modes—or real and imaginary parts of HPOD modes [29]—suggests that the Hilbert transform of a POD mode representing a propagating structure should closely match the POD mode corresponding to the $\pi/2$ phase-shift of the same structure.
@@ 3.1 Experimental Facility
Experiments were performed in a vertical water tunnel, the test section of which is 1.5 m long with a cross-section of 0.25 m $\times$ 0.25 m, shown in figure 1[7, 6, 11]. The settling chamber is connected to the top of the test section by a 16:1 contraction, which incorporates a honeycomb and screens to reduce the turbulence intensity and scale in the tunnel test section. Water is pumped from the plenum chamber below the test section up to the settling chamber such that the flow through the test section is downwards. The walls of the test section are made of 15 mm thick clear acrylic. A removable panel is located on one wall for access to the test section, and is set flush to the internal side of the wall to minimise asymmetry in the tunnel. Seeding and filtration of the tunnel is performed via an auxiliary circuit, which is isolated during experiments. A sting of diameter 9.3 mm is mounted to a crossbeam of thin aerofoil cross-sections above the test section. The sting tapers over the last 100 mm to a diameter of 4.65 mm and terminates in a threaded rod used to mount the sphere. A close-up view of the sphere mounting structure is shown in figure 1b.
A sphere with a diameter of $D=40$ mm was fabricated using a Phrozen Sonic Mighty 8k 3D printer, featuring a layer height of 35 µm and a planar resolution of 22 µm $\times$ 22 µm, as shown in figure 1c. The sphere has a shell thickness of 2.5 mm and an external boss at the top for mounting to the threaded rod on the end of the sting. The boss begins parallel to the sting and smoothly transitions to the sphere surface. The sphere is supported internally by a conical structure, which features slits to prevent uncured resin from being trapped in the cavities during printing.
Figure 1: (a) Vertical water tunnel facility, (b) close-up of the mounting structure and sphere, and (c) 3-D printed sphere design [11].
@@ 3.2 Experimental Method
Flow measurements in the wake of the sphere were conducted at a freestream velocity of $U_{\infty}=200$ mm/s, corresponding to a Reynolds number of $Re_{D}=7780$. Under these conditions, the background turbulence is 2%, and the mean convective time of the flow
@@
was 0.2 seconds [11]. The flow was seeded with 11 µm glass spheres, and single-exposed particle image pairs were recorded using a PCO Panda 26 DS camera with a 5120 $\times$ 5120 px array of 2.5 µm $\times$ 2.5 µm pixels. A Zeiss lens with a fixed 50 mm focal length was used to capture a 190 mm $\times$ 190 mm field of view, resulting in a spatial resolution of 37 µm/px. The field of view began just behind the trailing edge of the sphere, to avoid reflections off the surface, and extended to 5.2 diameters downstream of the sphere. The field of view was aligned with the centreline of the sphere, and extended to 2.3 diameters either side of the centreline in the transverse direction. The f-stop was set to 2.8, yielding a depth of field of 5.0 mm. The diffraction-limited minimum image diameter and expected particle image diameter for an 11 µm particle at these settings were 3.9 µm and 1.6 px, respectively.
Illumination was provided by 7 ns pulses from a pair of 400 mJ Spectra Physics Nd:YAG lasers. The beams were aligned collinearly and shaped into a sheet aligned to the tunnel centreline and the corresponding meridian of the sphere. The thickness of the laser sheet throughout the field of view was approximately 300 µm. Image pairs were captured with a separation of $\Delta t=1$ ms at an acquisition rate of 1 Hz ($\Delta T=1$ s), using a BeagleBone Black pulse generator [14] to synchronise the camera and laser pulses.
@@ 3.3 Data Processing
Mean background images of the first and second exposures were subtracted from the corresponding particle images. The instantaneous streamwise and transverse velocities, $u(x,y,t)$ and $v(x,y,t)$, were determined from the particle image pairs using two-component-two-dimensional (2C-2D) multigrid/multipass cross-correlation digital (MCCD) PIV [48, 36]. The final interrogation window size was 64 px $\times$ 32 px, with a grid spacing of 32 px $\times$ 16 px in the streamwise $x$ and transverse $y$ directions, respectively. This aspect ratio was selected to capture the higher streamwise velocity component and the higher velocity gradients in the transverse direction, while ensuring each interrogation window contained approximately 10 particles common to both images of each pair. The quality of the cross-correlation was increased through the application of an 8-point Hart filter. A maximum displacement limit of 20 px and normalised local median filter [47], with a threshold of 2 standard deviations, were applied to the raw displacement vectors. The validation criteria resulted in a rejection of 2.3% of vectors, which were replaced by an interpolation using a second-order polynomial fit of their 13 nearest neighbours [36]. A total of $N_{T}=12,600$ two-component velocity vector fields, each containing 50,000+ vectors, were determined. The instantaneous velocity fluctuations $u^{\prime}(x,y,t)$ and $v^{\prime}(x,y,t)$ were calculated by subtracting the mean velocity field, $\overline{u}(x,y)$ and $\overline{v}(x,y)$, from the instantaneous velocity field, $u(x,y,t)$ and $v(x,y,t)$.
The 95% confidence interval of the 2C-2D MCCDPIV algorithm used is 0.06 pixels [36], corresponding to a standard error of $\sigma_{\varepsilon_{u}}$ = 0.03 pixels. This translates to a standard error of $1.1\times 10^{-3}$ m/s in the instantaneous velocities. The uncertainty of the mean velocity arising from the random nature of the turbulent fluctuations can be estimated by including the mean streamwise Reynolds stress [3, 37]
@@
where $\sigma_{\varepsilon_{u}}$ is the standard error of the instantaneous streamwise velocity measurements. As the maxima of both $\overline{u^{\prime}u^{\prime}}$ and $\overline{v^{\prime}v^{\prime}}$ are of the order of $0.05U_{\infty}^{2}$ [11], a conservative estimate of the uncertainty in the mean velocity components is $4.0\times 10^{-4}$ m/s. The uncertainty in the fluctuations is thus
@@
Freestream velocity $U_{\infty}$ 200 mm/s -
Sphere diameter $D$ 40 mm -
Reynolds number $Re_{D}$ 7780 -
Convective time $t_{c}$ 0.2 s -
@@
Depth of field - 5.0 mm $0.125D$
Laser sheet thickness - 0.3 mm $0.0075D$
Diffraction-limited particle diameter - 3.9 µm
Spatial resolution SRSR 37 µm -
Grid spacing - 32 px $\times$ 16 px $0.03D\times 0.015D$
Interrogation window - 64 px $\times$ 32 px $0.06D\times 0.06D$
Field of view - 190 mm $\times$ 190 mm $4.7D\times 4.7D$
Number of 2C-2D velocity fields $N_{T}$ 12,600 -
Velocity vectors per field $N_{S}$  50,000+ -
Velocity vector validation rate - 97.7%
@@ 4.1 POD and HPOD Modes
POD was implemented by arranging the turbulent fluctuations into the matrix
@@
where $N_{S}$ denotes the number of velocity vectors per velocity field and $\boldsymbol{x}_{n}={(x,y)}_{n}$ represents the spatial coordinates of the $n^{th}$ vector, with fluctuating components $u^{\prime}$ and $v^{\prime}$ corresponding to the streamwise and transverse directions, respectively. Since $N_{S}>N_{T}$, the method of snapshots was employed to perform the POD. The two-instant correlation of $\boldsymbol{X}$ is expressed in matrix form as
@@
yields the eigenvectors $\boldsymbol{\Phi}$ and eigenvalues $\lambda$. The temporal coefficients are obtained from $\boldsymbol{\Phi}$
@@
and the spatial modes are given by projecting $\boldsymbol{X}$ onto $\boldsymbol{\Phi}$
@@
As the Fourier transform—and thus the Hilbert transform—assumes periodic signals, the matrix $\boldsymbol{X}$ was weighted in the streamwise direction using a Hann window [5] prior to performing HPOD to reduce spectral leakage in the modes. The weight at each streamwise index is defined as
$w_{\textrm{Hann}}(x)=\tanh\!\left(\alpha\sin^{2}(\pi n/N)\right),$ (35)
where $\alpha$ is a scaling parameter that controls the steepness of the weighting taper towards the domain edges, $n$ is the index of the streamwise location and $N$ is the total number of streamwise locations. A value of $\alpha=10$ was used to provide a relatively steep taper to zero at the domain edges while minimising the effect on the interior. The weighted analytic signal of $\boldsymbol{X}$
$\boldsymbol{X}^{a}=w_{\textrm{Hann}}(x)\boldsymbol{X}+i\mathcal{H}_{x}[w_{\textrm{Hann}}(x)\boldsymbol{X}],$ (36)
was determined using equations (18) and (19), and HPOD was performed by substituting $\boldsymbol{X}^{a}$ for $\boldsymbol{X}$ in equations (31) to (34). The resulting eigenvalues, temporal coefficients, and spatial modes are denoted by $\lambda^{a}_{k}$, $a^{a}_{k}(t)$, and $\psi^{a}_{k}(\boldsymbol{x})$, respectively.
The contribution of the leading HPOD modes to the TKE of $\boldsymbol{X}^{a}$ is significantly larger than that of the leading POD modes to the TKE of $\boldsymbol{X}$. The first, second, and third modes of the HPOD contribute 8.93%, 5.36%, and 4.05%, respectively, while those of the POD contribute 4.85%, 3.39%, and 2.55%, respectively. However, the difference in the contribution of the higher-order modes to the TKE is less pronounced, as shown in figure 2a. Consequently, the cumulative contribution to TKE of a given number of the HPOD modes is more than that of the equivalent number of POD modes, as shown in figure 2b and table 2. It should be noted that this is not a one-to-one comparison, as the total TKE of the analytic signal $\boldsymbol{X}^{a}$, 453,071 m2/s2, is greater than that of the data $\boldsymbol{X}$, 300,542 m2/s2. This difference arises from the imaginary component of $\boldsymbol{X}^{a}$, which is absent from $\boldsymbol{X}$.
@@
The first POD mode consists of strong fluctuations directly behind the sphere at $x/D<2.5$, with substantially weaker fluctuations present around $x/D=3.5$. The streamwise fluctuations, shown in figure 3a(i), are opposite in sign across the centreline, where they vanish, while the transverse fluctuations, shown in figure 3a(ii), alternate in sign along the streamwise direction and are maximal at the centreline. The second POD mode, the streamwise and transverse components of which are shown in figures 3b(i) and 3b(ii), respectively, appears similar to the first mode, with the structures shifted downstream. The structures present at $x/D<2.5$ in the first mode are now located at $1.5<x/D<3.5$, and new structures with opposite sign are present at $x/D<1.5$, suggesting that these modes represent a propagating structure of fluctuations with opposing sign. Although the third POD mode, shown in figure 3c, is asymmetric in the measured plane, it is similar in structure to the first and second modes, suggesting that the structure is not always planar-symmetric. The fourth and fifth POD modes, shown in figures 3d and 3e, respectively, both exhibit symmetric V-shaped structures in the streamwise fluctuations and antisymmetric structures in the transverse fluctuations. Although the streamwise and transverse components of the fifth POD mode appear more symmetric and antisymmetric, respectively, than those of the fourth mode, it still resembles the fourth mode shifted downstream.
@@
The sixth and eighth POD modes, shown in figures 3f and 3h, respectively, are similar in structure to the first and second POD modes but with a shorter wavelength. The seventh POD mode superficially resembles the fourth mode in its transverse component, shown in figure 3g(ii). However, the streamwise component, shown in figure 3g(i), does not exhibit the streamwise periodicity of the other leading modes. While this mode may represent an asymmetric aspect of the structure represented by the fourth and fifth modes, it is less clearly related to these modes than the third mode appears to be to the first and second modes. The ninth and tenth POD modes exhibit V-shaped structures in their streamwise components, shown in figures 3i(i) and 3j(i), respectively, similar to the fourth and fifth modes for $x/D>2.5$, which increase in wavelength further downstream. These modes differ in shape from the lower-order modes upstream of $x/D=2$, where the structures have an aspect ratio closer to unity than those in the higher-order modes. The transverse components of the ninth and tenth modes, shown in figures 3i(ii) and 3j(ii), respectively, resemble those of the fourth and fifth modes with a short wavelength. The wavelength of these modes also increases in the downstream direction.
@@
Figure 4: (i) Streamwise and (ii) transverse components of the first four HPOD modes. Real parts are shown in (a), (c), (e), and (g) and corresponding imaginary parts are shown in (b), (d), (f), and (h)
The first HPOD mode, the real and imaginary parts of which are shown in figure 4a and 4b, respectively, is qualitatively similar in structure to the first two POD modes, shown in figure 3a and 3b, respectively. The streamwise component of the first HPOD mode, the real and imaginary components of which are shown in figures 4a(i) and 4b(i), respectively, consists of alternating positive and negative regions in the streamwise direction and is antisymmetric across the centreline. The transverse component, the real and imaginary parts of which are shown in figures 4a(ii) and 4b(ii), respectively, consists of symmetric regions of alternating sign. This structure is very similar to that of the first two POD modes, except for the ends of the measurement domain, where the intensity is reduced by the Hann window. The small structures near the sphere at $x/D=1$, in the second POD mode, shown in figure 3b(i), are extended towards the centreline in the real component of the HPOD mode, similar to the downstream structures, due to the assumption of periodicity implicit in the Hilbert transform. In the imaginary part of the HPOD mode, shown in figure 4b(i), these structures are less pronounced as their intensity is reduced by the Hann window. The favouring of propagating structures by HPOD also results in an increase in the relative intensity of the downstream structures in figures 4a(i) and 4b(i) compared to those in figures 3a(i) and 3b(i).